# Simple Drop-in LoRA Conditioning on Attention Layers Will Improve Your Diffusion Model

### v1

@@ Abstract
Current state-of-the-art diffusion models employ U-Net architectures containing convolutional and (qkv) self-attention layers. The U-Net processes images while being conditioned on the time embedding input for each sampling step and the class or caption embedding input corresponding to the desired conditional generation. Such conditioning involves scale-and-shift operations to the convolutional layers but does not directly affect the attention layers. While these standard architectural choices are certainly effective, not conditioning the attention layers feels arbitrary and potentially suboptimal. In this work, we show that simply adding LoRA conditioning to the attention layers without changing or tuning the other parts of the U-Net architecture improves the image generation quality. For example, a drop-in addition of LoRA conditioning to EDM diffusion model yields FID scores of 1.91/1.75 for unconditional and class-conditional CIFAR-10 generation, improving upon the baseline of 1.97/1.79.
@@ 1 Introduction
In recent years, diffusion models have led to phenomenal advancements in image generation. Many cutting-edge diffusion models leverage U-Net architectures as their backbone, consisting of convolutional and (qkv) self-attention layers Dhariwal & Nichol (2021); Kim et al. (2023); Saharia et al. (2022); Rombach et al. (2022); Podell et al. (2024). In these models, the U-Net architecture-based score network is conditioned on the time, and/or, class, text embedding Ho & Salimans (2021) using scale-and-shift operations applied to the convolutional layers in the so-called residual blocks. Notably, however, the attention layers are not directly affected by the conditioning, and the rationale behind not extending conditioning to attention layers remains unclear. This gap suggests a need for in-depth studies searching for effective conditioning methods for attention layers and assessing their impact on performance.
Meanwhile, low-rank adaptation (LoRA) has become the standard approach for parameter-efficient fine-tuning of large language models (LLM) Hu et al. (2022). With LoRA, one trains low-rank updates that are added to frozen pre-trained dense weights in the attention layers of LLMs. The consistent effectiveness of LoRA for LLMs suggests that LoRA may be generally compatible with attention layers used in different architectures and for different tasks Chen et al. (2022); Pan et al. (2022); Lin et al. (2023); Gong et al. (2024).
In this work, we introduce a novel method for effectively conditioning the attention layers in the U-Net architectures of diffusion models by jointly training multiple LoRA adapters along with the base model. We call these LoRA adapters TimeLoRA and ClassLoRA for discrete-time settings, and Unified Compositional LoRA (UC-LoRA) for continuous signal-to-ratio (SNR) settings. Simply adding these LoRA adapters in a drop-in fashion without modifying or tuning the original model brings consistent enhancement in FID scores across several popular models applied to CIFAR-10, FFHQ 64x64, and ImageNet datasets. In particular, adding LoRA-conditioning to the EDM model Karras et al. (2022) yields improved FID scores of 1.75, 1.91, 2.31 for class-conditional CIFAR-10, unconditional CIFAR-10, and FFHQ 64x64 datasets, respectively, outperforming the baseline scores of 1.79, 1.97, 2.39. Moreover, we find that LoRA conditioning by itself is powerful enough to perform effectively. Our experiments show that only conditioning the attention layers using LoRA adapters (without the conditioning convolutional layers with scale-and-shift) achieves comparable FID scores compared to the baseline scale-and-shift conditioning (without LoRA).
@@
Our experiments show that using LoRA to condition time and class information on attention layers is effective across various models and datasets, including nano diffusion Lelarge et al. (2024), IDDPM Nichol & Dhariwal (2021), and EDM Karras et al. (2022) architectures using the MNIST Deng (2012), CIFAR-10 Krizhevsky et al. (2009), and FFHQ Karras et al. (2019) datasets.
Our main contributions are as follows. (i) We show that simple drop-in LoRA conditioning on the attention layers improves the image generation quality, as measured by lower FID scores, while incurring minimal ($\sim$10%) added memory and compute costs. (ii) We identify the problem of whether to and how to condition attention layers in diffusion models and provide the positive answer that attention layers should be conditioned and LoRA is an effective approach that outperforms the prior approaches of no conditioning or conditioning with adaLN Peebles & Xie (2023).
@@ 2.1 Diffusion models
Diffusion models Sohl-Dickstein et al. (2015); Song & Ermon (2019); Ho et al. (2020); Song et al. (2021b) generate images by iteratively removing noise from a noisy image. This denoising process is defined by the reverse process of the forward diffusion process: given data $x_{0}\sim q_{0}$, progressively inject noise to $x_{0}$ by
@@
A diffusion model is trained to approximate the score function $\nabla\log{p_{t}(x_{t})}$ with a score network $s_{\theta}$, which is often modeled with a U-Net architecture Ronneberger et al. (2015); Song & Ermon (2019). With $s_{\theta}\approx\nabla\log{p_{t}(x_{t})}$, the diffusion model approximates the reverse process as
@@
The above discrete-time description of diffusion models has a continuous-time counterpart based on the theory of stochastic differential equation (SDE) for the forward-corruption process and reversing it based on Anderson’s reverse-time SDE Anderson (1982) or a reverse-time ordinary differential equation (ODE) with equivalent marginal probabilities Song et al. (2021a). Higher-order integrators have been used to reduce the discretization errors in solving the differential equations Karras et al. (2022).
@@
The initial work of Song & Ermon (2019) first utilized the CNN-based U-Net architecture Ronneberger et al. (2015) as the architecture for the score network. Several improvements have been made by later works Ho et al. (2020); Nichol & Dhariwal (2021); Dhariwal & Nichol (2021); Hoogeboom et al. (2023) incorporating multi-head self-attention Vaswani et al. (2017), group normalization Wu & He (2018), and adaptive layer normalization (adaLN) Perez et al. (2018). Recently, several alternative architectures have been proposed. Jabri et al. (2023) proposed Recurrent Interface Network (RIN), which decouples the core computation and the dimension of the data for more scalable image generation. Peebles & Xie (2023); Bao et al. (2023); Gao et al. (2023); Hatamizadeh et al. (2023) investigated the effectiveness of transformer-based architectures Dosovitskiy et al. (2021) for diffusion models. Yan et al. (2023) utilized state space models Gu et al. (2022) in DiffuSSM to present an attention-free diffusion model architecture. In this work, we propose a conditioning method for attention layers and test it on several CNN-based U-Net architectures. Note that our proposed method is applicable to all diffusion models utilizing attention layers.
@@ 2.2 Low-rank adaptation
Using trainable adapters for specific tasks has been an effective approach for fine-tuning models in the realm of natural language processing (NLP) Houlsby et al. (2019); Pfeiffer et al. (2020). Low-rank adpatation (LoRA, Hu et al. (2022)) is a parameter-efficient fine-tuning method that updates a low-rank adapter: to fine-tune a pre-trained dense weight matrix $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, LoRA parameterizes the fine-tuning update $\Delta W$ with a low-rank factorization
@@
Although initially proposed for fine-tuning LLMs, LoRA is generally applicable to a wide range of other deep-learning modalities. Recent works used LoRA with diffusion models for various tasks including image generation Ryu (2023); Gu et al. (2023); Go et al. (2023), image editing Shi et al. (2023), continual learning Smith et al. (2023), and distillation Golnari (2023); Wang et al. (2023b). While all these works demonstrate the flexibility and efficacy of the LoRA architecture used for fine-tuning diffusion models, to the best of our knowledge, our work is the first attempt to use LoRA as part of the core U-Net for diffusion models for full training, not fine-tuning.
@@ 2.3 Conditioning the score network
For diffusion models to work properly, it is crucial that the score network $s_{\theta}$ is conditioned on appropriate side information. In the base formulation, the score function $\nabla_{x}p_{t}(x)$, which the score network $s_{\theta}$ learns, depends on the time $t$, so this $t$-dependence must be incorporated into the model via time conditioning. When class-labeled training data is available, class-conditional sampling requires class conditioning of the score network Ho & Salimans (2021). To take advantage of data augmentation and thereby avoid overfitting, EDM Karras et al. (2022) utilizes augmentation conditioning Jun et al. (2020), where the model is conditioned on the data augmentation information such as the degree of image rotation or blurring. Similarly, SDXL Podell et al. (2024) uses micro-conditioning, where the network is conditioned on image resolution or cropping information.
Finally, text-to-image diffusion models Saharia et al. (2022); Ramesh et al. (2022); Rombach et al. (2022); Podell et al. (2024) use text conditioning, which conditions the score network with caption embeddings so that the model generates images aligned with the text description.
@@
There is a line of research proposing transformer-based architectures (without convolutions) for diffusion models, and these work do propose methods for conditioning attention layers. For instance, DiT Peebles & Xie (2023) conditioned attention layers using adaLN and DiffiT Hatamizadeh et al. (2023) introduced time-dependent multi-head self-attention (TMSA), which can be viewed as scale-and-shift conditioning applied to attention layers. Although such transformer-based architectures have shown to be effective, whether conditioning the attention layers with adaLN or scale-and-shift is optimal was not investigated. In Section 5.5 of this work, we compare our proposed LoRA conditioning on attention layers with the prior adaLN conditioning on attention layers, and show that LoRA is the more effective mechanism for conditioning attention layers.
@@
Multi-task learning Caruana (1997) is a framework where a single model is trained on multiple related tasks simultaneously, leveraging shared representations between the tasks. If one views the denoising tasks for different timesteps (or SNR) of diffusion models as related but different tasks, the training of diffusion models can be interpreted as an instance of the multi-task learning. Following the use of trainable lightweight adapters for Mixture-of-Expert (MoE) Jacobs et al. (1991); Ma et al. (2018), several works have utilized LoRA as the expert adapter for the multi-task learning Caccia et al. (2023); Wang et al. (2023a; 2024); Zadouri et al. (2024). Similarly, MORRIS Audibert et al. (2023) and LoRAHub Huang et al. (2023) proposed using the weighted sum of multiple LoRA adapters to effectively tackle general tasks. In this work, we took inspiration from theses works by using a composition of LoRA adapters to condition diffusion models.
@@ 3 Discrete-time LoRA conditioning
Diffusion models such as DDPM Ho et al. (2020) and IDDPM Nichol & Dhariwal (2021) have a predetermined number of discrete timesteps $t=1,2,\dots,T$ used for both training and sampling. We refer to this setting as the discrete-time setting.
@@ 3.2 ClassLoRA
for $c=1,\dots,C$. In other words, each LoRA $(A^{\prime}_{c},B^{\prime}_{c})$ handles a single class $c$. When $C$ is large, such as in the case of ImageNet1k, one may consider using a compositional version of ClassLoRA.
@@ 4 Continuous-SNR LoRA conditioning
Motivated by (Kingma et al., 2021), some recent models such as EDM Karras et al. (2022) consider parameterizing the score function as a function of noise or signal-to-noise ratio (SNR) level instead of time. In particular, EDM Karras et al. (2022) considers the probability flow ODE
@@
The main distinction between Sections 3 and 4 is in the discrete vs. continuous parameterization, since continuous-time and continuous-SNR parameterizations of score functions are equivalent. We choose to consider continuous-SNR (instead of continuous-time) parameterizations for the sake of consistency with the EDM model Karras et al. (2022).
@@ 5.1 Experimental Setup
We implement LoRA conditioning on three different diffusion models: nano diffusion Lelarge et al. (2024), IDDPM Nichol & Dhariwal (2021), and EDM-vp Karras et al. (2022). With nano diffusion, we conduct a proof-of-concept experiment. With IDDPM, we test TimeLoRA and ClassLoRA for the discrete-time setting, and with EDM, we test UC-LoRA for the continuous-SNR setting.
@@
We use the standard LoRA initialization as in the original LoRA paper Hu et al. (2022): for the LoRA matrices ($A$, $B$) with rank $r$, $A$ is initialized as $A_{ij}\sim\mathcal{N}(0,1/r)$ and $B$ as the zero matrix. Following Ryu (2023), we set the rank of each basis LoRA to $4$. For TimeLoRA and ClassLoRA, we use $11$ and $10$ LoRA bases, and for UC-LoRA we use $18$ and $20$ LoRA bases for CIFAR-10 and FFHQ.
@@ 5.3 Main quantitative results
To evaluate the effectiveness of the drop-in addition of LoRA conditioning to the attention layers, we implement TimeLoRA and ClassLoRA to IDDPM and UC-LoRA to EDM, both with the conventional scale-and-shift conditioning on the convolutional layers unchanged. We train IDDPM with CIFAR-10, ImageNet64 and EDM with CIFAR-10, FFHQ64. As reported in Table 1, the addition of LoRA conditioning to the attention layers consistently improves the image generation quality as measured by FID scores Heusel et al. (2017) across different diffusion models and datasets with only ($\sim$10%) addition of the parameter counts. Note these improvements are achieved without tuning any hyperparameters of the base model components.

### v3

@@ Abstract
Current state-of-the-art diffusion models employ U-Net architectures containing convolutional and (qkv) self-attention layers. The U-Net processes images while being conditioned on the time embedding input for each sampling step and the class or caption embedding input corresponding to the desired conditional generation. Such conditioning involves scale-and-shift operations to the convolutional layers but does not directly affect the attention layers. While these standard architectural choices are certainly effective, not conditioning the attention layers feels arbitrary and potentially suboptimal. In this work, we show that simply adding LoRA conditioning to the attention layers without changing or tuning the other parts of the U-Net architecture improves the image generation quality. For example, a drop-in addition of LoRA conditioning to EDM diffusion model yields FID scores of 1.91/1.75 for unconditional and class-conditional CIFAR-10 generation, improving upon the baseline of 1.97/1.79. Code is available at https://github.com/lthilnklover/diffusion_lora.
@@ 1 Introduction
In recent years, diffusion models have led to phenomenal advancements in image generation. Many cutting-edge diffusion models leverage U-Net architectures as their backbone, consisting of convolutional and (qkv) self-attention layers (Dhariwal & Nichol, 2021; Kim et al., 2023; Saharia et al., 2022; Rombach et al., 2022; Podell et al., 2024). In these models, the U-Net architecture-based score network is conditioned on the time, and/or, class, text embedding (Ho & Salimans, 2021) using scale-and-shift operations applied to the convolutional layers in the so-called residual blocks. Notably, however, the attention layers are not directly affected by the conditioning, and the rationale behind not extending conditioning to attention layers remains unclear. This gap suggests a need for in-depth studies searching for effective conditioning methods for attention layers and assessing their impact on performance.
Meanwhile, low-rank adaptation (LoRA) has become the standard approach for parameter-efficient fine-tuning of large language models (LLM) (Hu et al., 2022). With LoRA, one trains low-rank updates that are added to frozen pre-trained dense weights in the attention layers of LLMs. The consistent effectiveness of LoRA for LLMs suggests that LoRA may be generally compatible with attention layers used in different architectures and for different tasks (Chen et al., 2022; Pan et al., 2022; Lin et al., 2023; Gong et al., 2024).
In this work, we introduce a novel method for effectively conditioning the attention layers in the U-Net architectures of diffusion models by jointly training multiple LoRA adapters along with the base model. We call these LoRA adapters TimeLoRA and ClassLoRA for discrete-time settings, and Unified Compositional LoRA (UC-LoRA) for continuous signal-to-ratio (SNR) settings. Simply adding these LoRA adapters in a drop-in fashion without modifying or tuning the configurations of original model brings consistent enhancement in FID scores across several popular models applied to CIFAR-10, FFHQ 64x64, and ImageNet datasets. In particular, adding LoRA-conditioning to the EDM model (Karras et al., 2022) yields improved FID scores of 1.75, 1.91, 2.28 for class-conditional CIFAR-10, unconditional CIFAR-10, and FFHQ 64x64 datasets, respectively, outperforming the baseline scores of 1.79, 1.97, 2.39. Moreover, we find that LoRA conditioning by itself is powerful enough to perform effectively. Our experiments show that only conditioning the attention layers using LoRA adapters (without the conditioning convolutional layers with scale-and-shift) achieves comparable FID scores compared to the baseline scale-and-shift conditioning (without LoRA).
@@
Our experiments show that using LoRA to condition time and class information on attention layers is effective across various models and datasets, including nano diffusion (Lelarge et al., 2024), IDDPM (Nichol & Dhariwal, 2021), and EDM (Karras et al., 2022) architectures using the MNIST (Deng, 2012), CIFAR-10 (Krizhevsky et al., 2009), and FFHQ (Karras et al., 2019) datasets.
Our main contributions are as follows. (i) We show that simple drop-in LoRA conditioning on the attention layers improves the image generation quality, as measured by lower FID scores, while incurring minimal ($\sim$10%) added memory and compute costs. (ii) We identify the problem of whether to and how to condition attention layers in diffusion models and provide the positive answer that attention layers should be conditioned and LoRA is an effective approach that outperforms the prior approaches of no conditioning or conditioning with adaLN (Peebles & Xie, 2023).
@@ 2.1 Diffusion models
Diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Song et al., 2021b) generate images by iteratively removing noise from a noisy image. This denoising process is defined by the reverse process of the forward diffusion process: given data $x_{0}\sim q_{0}$, progressively inject noise to $x_{0}$ by
@@
A diffusion model is trained to approximate the score function $\nabla\log{p_{t}(x_{t})}$ with a score network $s_{\theta}$, which is often modeled with a U-Net architecture (Ronneberger et al., 2015; Song & Ermon, 2019). With $s_{\theta}\approx\nabla\log{p_{t}(x_{t})}$, the diffusion model approximates the reverse process as
@@
The above discrete-time description of diffusion models has a continuous-time counterpart based on the theory of stochastic differential equation (SDE) for the forward-corruption process and reversing it based on Anderson’s reverse-time SDE (Anderson, 1982) or a reverse-time ordinary differential equation (ODE) with equivalent marginal probabilities (Song et al., 2021a). Higher-order integrators have been used to reduce the discretization errors in solving the differential equations (Karras et al., 2022).
@@
The initial work of Song & Ermon (2019) first utilized the CNN-based U-Net architecture (Ronneberger et al., 2015) as the architecture for the score network. Several improvements have been made by later works (Ho et al., 2020; Nichol & Dhariwal, 2021; Dhariwal & Nichol, 2021; Hoogeboom et al., 2023) incorporating multi-head self-attention (Vaswani et al., 2017), group normalization (Wu & He, 2018), and adaptive layer normalization (adaLN) (Perez et al., 2018). Recently, several alternative architectures have been proposed. Jabri et al. (2023) proposed Recurrent Interface Network (RIN), which decouples the core computation and the dimension of the data for more scalable image generation. Peebles & Xie (2023); Bao et al. (2023); Gao et al. (2023); Hatamizadeh et al. (2023) investigated the effectiveness of transformer-based architectures (Dosovitskiy et al., 2021) for diffusion models. Yan et al. (2023) utilized state space models (Gu et al., 2022) in DiffuSSM to present an attention-free diffusion model architecture. In this work, we propose a conditioning method for attention layers and test it on several CNN-based U-Net architectures. Note that our proposed method is applicable to all diffusion models utilizing attention layers.
@@ 2.2 Low-rank adaptation
Using trainable adapters for specific tasks has been an effective approach for fine-tuning models in the realm of natural language processing (NLP) (Houlsby et al., 2019; Pfeiffer et al., 2020). Low-rank adpatation (LoRA, Hu et al. (2022)) is a parameter-efficient fine-tuning method that updates a low-rank adapter: to fine-tune a pre-trained dense weight matrix $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, LoRA parameterizes the fine-tuning update $\Delta W$ with a low-rank factorization
@@
Although initially proposed for fine-tuning LLMs, LoRA is generally applicable to a wide range of other deep-learning modalities. Recent works used LoRA with diffusion models for various tasks including image generation (Ryu, 2023; Gu et al., 2023; Go et al., 2023), image editing (Shi et al., 2023), continual learning (Smith et al., 2023), and distillation (Golnari, 2023; Wang et al., 2023b). While all these works demonstrate the flexibility and efficacy of the LoRA architecture used for fine-tuning diffusion models, to the best of our knowledge, our work is the first attempt to use LoRA as part of the core U-Net for diffusion models for full training, not fine-tuning.
@@ 2.3 Conditioning the score network
For diffusion models to work properly, it is crucial that the score network $s_{\theta}$ is conditioned on appropriate side information. In the base formulation, the score function $\nabla_{x}p_{t}(x)$, which the score network $s_{\theta}$ learns, depends on the time $t$, so this $t$-dependence must be incorporated into the model via time conditioning. When class-labeled training data is available, class-conditional sampling requires class conditioning of the score network (Ho & Salimans, 2021). To take advantage of data augmentation and thereby avoid overfitting, EDM (Karras et al., 2022) utilizes augmentation conditioning (Jun et al., 2020), where the model is conditioned on the data augmentation information such as the degree of image rotation or blurring. Similarly, SDXL (Podell et al., 2024) uses micro-conditioning, where the network is conditioned on image resolution or cropping information. Finally, text-to-image diffusion models (Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Podell et al., 2024) use text conditioning, which conditions the score network with caption embeddings so that the model generates images aligned with the text description.
@@
There is a line of research proposing transformer-based architectures (without convolutions) for diffusion models, and these work do propose methods for conditioning attention layers. For instance, DiT (Peebles & Xie, 2023) conditioned attention layers using adaLN and DiffiT (Hatamizadeh et al., 2023) introduced time-dependent multi-head self-attention (TMSA), which can be viewed as scale-and-shift conditioning applied to attention layers. Although such transformer-based architectures have shown to be effective, whether conditioning the attention layers with adaLN or scale-and-shift is optimal was not investigated. In Section 5.5 of this work, we compare our proposed LoRA conditioning on attention layers with the prior adaLN conditioning on attention layers, and show that LoRA is the more effective mechanism for conditioning attention layers.
@@
Multi-task learning (Caruana, 1997) is a framework where a single model is trained on multiple related tasks simultaneously, leveraging shared representations between the tasks. If one views the denoising tasks for different timesteps (or SNR) of diffusion models as related but different tasks, the training of diffusion models can be interpreted as an instance of the multi-task learning. Following the use of trainable lightweight adapters for Mixture-of-Expert (MoE) (Jacobs et al., 1991; Ma et al., 2018), several works have utilized LoRA as the expert adapter for the multi-task learning (Caccia et al., 2023; Wang et al., 2023a; 2024; Zadouri et al., 2024). Similarly, MORRIS (Audibert et al., 2023) and LoRAHub (Huang et al., 2023) proposed using the weighted sum of multiple LoRA adapters to effectively tackle general tasks. In this work, we took inspiration from theses works by using a composition of LoRA adapters to condition diffusion models.
@@ 3 Discrete-time LoRA conditioning
Diffusion models such as DDPM (Ho et al., 2020) and IDDPM (Nichol & Dhariwal, 2021) have a predetermined number of discrete timesteps $t=1,2,\dots,T$ used for both training and sampling. We refer to this setting as the discrete-time setting.
@@ 3.2 ClassLoRA
for $c=1,\dots,C$. In other words, each LoRA $(A^{\prime}_{c},B^{\prime}_{c})$ handles a single class $c$. When $C$ is large, such as in the case of ImageNet1k, one may consider using a compositional version of ClassLoRA (See Appendix B).
@@ 4 Continuous-SNR LoRA conditioning
Motivated by (Kingma et al., 2021), some recent models such as EDM (Karras et al., 2022) consider parameterizing the score function as a function of noise or signal-to-noise ratio (SNR) level instead of time. In particular, EDM (Karras et al., 2022) considers the probability flow ODE
@@
The main distinction between Sections 3 and 4 is in the discrete vs. continuous parameterization, since continuous-time and continuous-SNR parameterizations of score functions are equivalent. We choose to consider continuous-SNR (instead of continuous-time) parameterizations for the sake of consistency with the EDM model (Karras et al., 2022).
@@ 5.1 Experimental Setup
We implement LoRA conditioning on three different diffusion models: nano diffusion (Lelarge et al., 2024), IDDPM (Nichol & Dhariwal, 2021), and EDM-vp (Karras et al., 2022). With nano diffusion, we conduct a proof-of-concept experiment. With IDDPM, we test TimeLoRA and ClassLoRA for the discrete-time setting, and with EDM, we test UC-LoRA for the continuous-SNR setting.
@@
We use the standard LoRA initialization as in the original LoRA paper (Hu et al., 2022): for the LoRA matrices ($A$, $B$) with rank $r$, $A$ is initialized as $A_{ij}\sim\mathcal{N}(0,1/r)$ and $B$ as the zero matrix. Following Ryu (2023), we set the rank of each basis LoRA to $4$. For TimeLoRA and ClassLoRA, we use $11$ and $10$ LoRA bases, and for UC-LoRA we use $18$ and $20$ LoRA bases for CIFAR-10 and FFHQ.
@@ 5.3 Main quantitative results
To evaluate the effectiveness of the drop-in addition of LoRA conditioning to the attention layers, we implement TimeLoRA and ClassLoRA to IDDPM and UC-LoRA to EDM, both with the conventional scale-and-shift conditioning on the convolutional layers unchanged. We train IDDPM with CIFAR-10, ImageNet64 and EDM with CIFAR-10, FFHQ64. As reported in Table 1, the addition of LoRA conditioning to the attention layers consistently improves the image generation quality as measured by FID scores (Heusel et al., 2017) across different diffusion models and datasets with only ($\sim$10%) addition of the parameter counts. Note these improvements are achieved without tuning any hyperparameters of the base model components.